# Exploring Learning Complexity for Efficient Downstream Dataset Pruning

### v2

@@ Abstract
The ever-increasing fine-tuning cost of large-scale pre-trained models gives rise to the importance of dataset pruning, which aims to reduce dataset size while maintaining task performance. However, existing dataset pruning methods require training on the entire dataset, which is impractical for large-scale pre-trained models. In this paper, we propose a straightforward, novel, and training-free hardness score named Distorting-based Learning Complexity (DLC), to identify informative images and instructions from the downstream dataset efficiently. Our method is motivated by the observation that easy samples learned faster can also be learned with fewer parameters. Specifically, we define the Learning Complexity to quantify sample hardness and utilize a lightweight weights masking process for fast estimation, instead of the costly SGD optimization. Based on DLC, we further design a flexible under-sampling with randomness (dubbed FlexRand), replacing the top-K strategy, to alleviate the severe subset distribution shift. Extensive experiments with downstream image and instructions dataset pruning benchmarks demonstrate the effectiveness and efficiency of the proposed approach. In the images pruning benchmark, DLC significantly reduces the pruning time by 35$\times$ while establishing state-of-the-art performance with FlexRand.
@@ 1 Introduction
In this work, we propose a training-free hardness score, implementing the Learning Complexity (Definition 2) by a lightweight pre-training parameters masking process. From the hardness perspective, the above learning complexity aims to quantify sample importance via calculating the classification loss integration over a model sequence with ever-increasing classification performance, which is defined as the Learning Path (Definition 1). The proposed scoring function, Distorting-based Learning Complexity (dubbed DLC), is an efficient learning path implementation, by the observation that distorting (masking) the pre-training weights can construct a learning path without training or fine-tuning. Specifically, we mask pre-training parameters multiple times with random ratios, and the learning complexity, i.e., loss integration, can be efficiently approximated by the Monte Carlo method (Caflisch, 1998). As a result, DLC can distinguish easy samples from hard ones. To further verify the effectiveness of DLC, we empirically demonstrate a high ranking correlation between DLC and the common optimization-based learning complexity. Moreover, we extend DLC to quantify the instruction hardness for efficiently fine-tuning the large language models (LLMs).
To identify informative samples based on DLC, we design a flexible under-sampling strategy with randomness, named FlexRand, replacing the common top-K strategy. For the multinomial distribution of under-sampling, FlexRand aims to flexibly adjust sampling preference at different data regimes. Moreover, the subset from FlexRand should not be significantly shifted away from the original data distribution. Specifically, we randomly select the same number of samples from the easy and hard intervals respectively, whose lengths are flexibly tuned with a splitting hyper-parameter $\gamma$ to accommodate different data regimes. In this way, we have the following two advantages: (a): FlexRand can adapt to different data regimes; (b): FlexRand avoids severe distribution shift.
@@ 3 Method
To design an efficient scoring function, we first define the Learning Complexity to quantify sample importance from the hardness perspective. Specifically, we calculate the integration of classification loss for a sample predicted by an improving model sequence, which is defined as:
@@
Figure 2: Ranking correlation between the loss integration over the optimization and masking process. We fine-tune the pre-trained ResNet-18 on the five downstream datasets for 50 epochs and present the CXRB10 (Nodule) results here due to the space limit (see Figure 7 for full results). The easy (hard) sample corresponds to the data point with a minimum (maximum) loss integration over the optimization. (a): Loss trends with the number of parameters, i.e., model capacity. We produce models with different capacities via masking {2%, …, 98%} weights of the pre-trained ResNet-18. Average category representation acts as the prototype for classification and loss integration. (b): Loss trends with the optimization time. (c): High ranking correlation coefficient with $\rho=\mathbf{0.54}$.
@@ 3.1 DLC: Distorted-based Learning Complexity
DLC is motivated by the observation that increasing the number of parameters leads to improved generalization (Valle-Perez et al., 2019; Neyshabur et al., 2018; 2017). Although the pre-trained models have not been fine-tuned on the downstream datasets, the expected risk (test error) on the target task still decreases with the increasing model capacity as shown in Figure 2(a). Formally, when the capacity of pre-trained model $f_{\bm{\hat{\theta}}}$ at step $t+1$ is larger than that at step $t$, we have
$\mathcal{R}_{\mathcal{L}}(f_{\bm{\hat{\theta}}(t+1)})<\mathcal{R}_{\mathcal{L}}(f_{\bm{\hat{\theta}}(t)})$
Therefore, the sequence of pre-trained models with different capacities is viable and can be constructed efficiently without fine-tuning. In detail, we directly produce the classifiers with different capacities by masking the initial pre-training weights with ratios ranging from $\{2\%,4\%,...,98\%,100\%\}$. Note that parameters with small L1 norm are masked first (Han et al., 2015). To calculate the integration of classification loss, i.e., Learning Complexity, the average category feature acts as the class prototype, and we further adopt the Monte Carlo method to approximate the loss integration (Caflisch, 1998) for efficiency, instead of enumerating all models with different capacities:
@@
To empirically verify the asymptotic convergence and decreasing generalization error, we visualize the test loss trend in Figure 2(b). Obviously, $f_{\bm{\theta}(0)}$ gradually converges to $f_{\bm{{\hat{\theta}}}}$ with better generalization performance. Following the Definition 1, the optimization trajectory of SGD during the fine-tuning is a viable learning path, and we calculate the integration of classification loss over the fine-tuning process as the optimization-based learning complexity.
@@ 3.2 FlexRand: Flexible Under-sampling with Randomness
From our previous analysis, we show that DLC can discriminate the hard samples from the easy. To identify informative samples with DLC, we further design a flexible under-sampling strategy with randomness, named FlexRand, replacing the common top-K strategy. For the multinomial distribution of under-sampling, FlexRand aims to flexibly adjust sampling preference at different data regimes. On the other hand, the subset from FlexRand should not be significantly shifted away from the original data distribution. To achieve these objectives, we randomly select the same number of samples from the easy and hard intervals respectively, whose lengths are flexibly tuned with a splitting hyper-parameter $\gamma$ to accommodate different data regimes. The sampling probability for $(\bm{x},y)$ can be formulated as follows:
@@
In a data-poor regime, easy samples are more informative, and the opposite holds when more samples can be preserved for fine-tuning (Sorscher et al., 2022). However, the top-K (Easy or Hard) strategy, denoted as T-E or T-H, fixes sampling preference across different data regimes and shows inferior downstream performance compared to the random as shown in Figure 4(b) & 4(d). On the contrary, FlexRand can flexibly adjust sampling preference with a splitting hyper-parameter $\gamma$ to accommodate different data regimes.
@@
Figure 4: Harmful distribution shift from the top-K under-sampling strategies. We conduct downstream dataset pruning with four different under-sampling strategies according to the optimization-based learning complexity and DLC. (a): Distribution shift comparison of four subsets for optimization-based learning complexity. (b): Downstream performance comparison of four subsets for optimization-based learning complexity. (c): Distribution shift comparison of four subsets for DLC. (d): Downstream performance comparison of four subsets for DLC.
@@ 3.3 Extension to Instruction Dataset Pruning
Recently, large language models (LLMs) (Touvron et al., 2023a; b; Jiang et al., 2024) driven by the PT-FT paradigm have shown incredible capabilities across a wide range of language tasks. Despite the availability of pre-trained weights, the ever-increasing number of parameters still incur heavy computation overhead for the fine-tuning, which involves instruction-tuning (Ouyang et al., 2022) and reinforcement learning with human feedback (RLHF) (Knox & Stone, 2011) to align the pre-trained LLMs with human preferences. On the other hand, recent works (Zhou et al., 2023; Cao et al., 2023) indicate that almost all knowledge in LLMs is learned during pre-training, and only limited instruction tuning data is necessary to teach models to produce high-quality output. Therefore, we further extend the learning complexity to prune instruction datasets for efficient large language models fine-tuning (Dubey et al., 2024; Team et al., 2024). Specifically, we construct the learning path by masking different numbers of weights in the pre-trained large language models (LLMs), which are sorted by L1 norm in ascending order. Due to the prediction difference in image classification and text generation, we replace the original loss function $\mathcal{L}(f(\bm{x};\bm{\theta},y)$ with:
@@ 4 Experiments
In this section, we present the diverse downstream image and instruction datasets pruning benchmarks and empirically validate the effectiveness and efficiency of the proposed method. Moreover, we perform ablation studies to understand better how different components and hyper-parameters influence the performance. The code is available in the supplementary material.
Table 1: Average classification accuracy over 5 diverse downstream datasets and pruning ratios with varying model architectures. Bold numbers are the optimal results, and underline numbers are suboptimal results. Detailed results can be found in Appendix B.2.2.
@@
Figure 5: Trends of average accuracy gap over the random method at different pruning ratios. Detailed results can be found in Appendix B.2.2.
@@ 4.1.1 Downstream Image Dataset Pruning Benchmark
For a comprehensive evaluation, we adopt pre-training encoders with different depths, architectures, and pre-training paradigms. In detail, the pre-trained encoders consist of ResNet-18, ResNet-50 (He et al., 2016), ViT-Small, and ViT-Base (Dosovitskiy, 2020). Those models are pretrained on the canonical ImageNet-1K (Deng et al., 2009) dataset via fully and weakly supervised learning (Yalniz et al., 2019) respectively.
@@
Generally, the downstream dataset is domain-specific and different from the pre-training dataset ImageNet-1K. Therefore, we choose diverse downstream datasets from 5 domains (Islam et al., 2021) to construct the large-scale benchmark, including CXRB10222CXRB10 is created by selecting 10 balanced classes from the ChestX-ray14 (Borghesi & Maroldi, 2020)., DeepWeeds (Olsen et al., 2019), DTD (Cimpoi et al., 2014), FGVCAircraft (Maji et al., 2013), and Sketch (Eitz et al., 2012). For hyper-parameter tuning, we split 20% as the validation set.
@@ 4.1.2 Downstream Instruction Dataset Pruning Benchmark
To align the based pre-trained models, we adopt 2 different instruction fine-tuning datasets: Alpaca Cleaned (Taori et al., 2023) and Dolly & HH-RLHF [^3] https://huggingface.co/datasets/mosaicml/dolly_hhrlhf.
More setup details are presented in Appendix B.1.
@@ 4.2.1 Fine-tuning for Image Classification
Our method achieves superior accuracy with different architectures.
@@
Our method consistently outperforms the random method with diverse downstream setups.
For efficacy verification of our method on diverse downstream setups, we compare the accuracy gap trends over the random strategy for each fine-tuning dataset. As shown in Figure 5, the proposed method shows consistent improvement for each downstream dataset at different data regimes, indicating the efficacy of the flexible under-sampling strategy with randomness. Overall, the advantage is coherent and more evident when fine-tuning with fewer samples. For example, our method outperforms the random baseline by 2.6% when removing 90% samples from the Sketch dataset. Despite the randomness in the custom under-sampling, our method demonstrates stability with smaller performance variance. However, the compared methods show inferior performance with fluctuating trends, which reveals the drawbacks of the fixed sampling principle and the challenge of our diverse downstream dataset pruning benchmark.
Our method significantly reduces the time cost of downstream dataset pruning.
In addition to the accuracy evaluated on different models and downstream setups, we also record the time cost of each baseline, including hyper-parameter tuning associated with the scoring estimation and under-sampling. In Table 1, time results averaged over 5 downstream datasets are presented. Benefiting from the training-free scoring function and efficient hyper-parameter tuning, the proposed method significantly reduces the pruning cost by 35$\times$ compared with the one-shot fine-tuning baselines while achieving state-of-the-art downstream performance. Notably the acceleration is more remarkable with ever-increasing model capacities. Therefore, our method gives rise to the efficient fine-tuning of the over-parameterized pre-trained models.
@@ 4.2.2 Fine-tuning LLMs for Text Classification
Figure 6: Results for different ablation studies. (a): Ablation on the learning path size. (b): Ablation on the masking strategies. (c): Effectiveness of FlexRand. (d): Ablation on the splitting $\gamma$.
@@ 4.3 Ablation Studies
To better understand why our method can obtain superior efficiency and efficacy on the downstream dataset pruning task, we perform extensive ablation studies on the learning complexity metric, under-sampling strategy, and corresponding hyper-parameters. As for the setups, we fine-tune the fully pre-trained ResNet-18 on 5 downstream datasets with 9 pruning ratios with default parameters.
@@
In Figure 6(a), we investigate the effect of the learning path size on the ranking correlation and downstream accuracy. With an ever-increasing learning path from 1 to 50, class-averaged Spearman correlation coefficients between the optimization-based learning complexity and masking-based are presented. We can observe that the ranking correlation between the above two scoring functions is significant and saturated at size 5. When the size is in the interval 5$\sim$10, the ranking correlation no longer changes, and the hyper-parameter tuning method stops searching. As a result, the proposed method estimates the learning complexity efficiently and establishes superior classification performance.
@@
To examine the effects of interval splitting, we set the $\gamma$ by traversing {0.05, 0.10, …, 0.95}, and the results for different downstream datasets are presented separately in Figure 6(d). We can find that small splitting values ($\gamma<0.5$) generally deliver better performance in data-poor regimes. Therefore, our method outperforms the random baseline. While $\gamma$ from tuning with the linear classifier does not always achieve the optimal, it significantly reduces the cost of tuning hyper-parameter and still establishes superior performance.

### v3

@@ Abstract
The ever-increasing fine-tuning cost of large-scale pre-trained models gives rise to the importance of dataset pruning, which aims to reduce dataset size while maintaining task performance. However, existing dataset pruning methods require training on the entire dataset, which is impractical for large-scale pre-trained models. In this paper, we propose a straightforward, novel, and training-free hardness score named Distorting-based Learning Complexity (DLC), to identify informative images and instructions from the downstream dataset efficiently. Our method is motivated by the observation that easy samples learned faster can also be learned with fewer parameters. Specifically, we define the Learning Complexity to quantify sample hardness and utilize a lightweight weights masking process for fast estimation, instead of the costly SGD optimization. Based on DLC, we further design a flexible under-sampling strategy with randomness (dubbed FlexRand), replacing the top-K strategy, to alleviate the severe subset distribution shift. Extensive experiments with downstream image and instructions dataset pruning benchmarks demonstrate the effectiveness and efficiency of the proposed approach. In the images pruning benchmark, DLC significantly reduces the pruning time by 35$\times$ while establishing state-of-the-art performance with FlexRand.
@@ 1 Introduction
In this work, we propose a training-free hardness score, implementing the Learning Complexity (Definition 2) by a lightweight pre-training parameters masking process. From the hardness perspective, the above learning complexity aims to quantify sample importance via calculating the classification loss integral over a model sequence with ever-increasing classification performance, which is defined as the Learning Path (Definition 1). The proposed scoring function, Distorting-based Learning Complexity (dubbed DLC), is an efficient learning complexity implementation, by the observation that distorting (masking) the pre-training weights can construct a learning path without training or fine-tuning. Specifically, we mask pre-training weights multiple times with random ratios, and the learning complexity can be efficiently approximated by the Monte Carlo method (Caflisch, 1998). As a result, DLC can distinguish easy samples from hard ones. To further verify the effectiveness of DLC, we empirically demonstrate a high ranking correlation between DLC and the common optimization-based learning complexity. Moreover, we extend DLC to quantify the instruction hardness for efficiently fine-tuning the large language models (LLMs).
To identify informative samples based on DLC, we design a flexible under-sampling strategy with randomness, named FlexRand, replacing the common top-K strategy. For the multinomial distribution of under-sampling, FlexRand aims to flexibly adjust sampling preference at different data regimes. Moreover, the subset from FlexRand should not be significantly shifted away from the original data distribution. Specifically, we randomly select the same number of samples from the easy and hard intervals respectively, whose lengths are flexibly tuned with a splitting hyperparameter $\gamma$ to accommodate different data regimes. In this way, we have the following two advantages: (a): FlexRand can adapt to different data regimes; (b): FlexRand avoids severe distribution shift.
@@ 3 Method
To design an efficient scoring function, we first define the Learning Complexity to quantify sample importance from the hardness perspective. Specifically, we calculate the integral of classification loss for a sample predicted by an improving model sequence, which is defined as:
@@
Figure 2: Ranking correlation between the loss integral over the optimization and masking process. We fine-tune the pre-trained ResNet-18 on the five downstream datasets for 50 epochs and present the CXRB10 (Nodule) results here due to the space limit (see Figure 7 for full results). The easy (hard) sample corresponds to the data point with a minimum (maximum) loss integral over the optimization. We utilize Max-Min normalization to scale the loss to the range of (0, 1). (a): Loss trends with the number of parameters, i.e., model capacity. We produce models with different capacities via masking {2%, …, 98%} weights of the pre-trained ResNet-18. Average category representation acts as the prototype for classification and loss integral. (b): Loss trends with the optimization time. (c): High ranking correlation coefficient with $\rho=\mathbf{0.54}$.
@@ 3.1 DLC: Distorted-based Learning Complexity
DLC is motivated by the observation that increasing the number of parameters leads to improved generalization (Valle-Perez et al., 2019; Neyshabur et al., 2018; 2017). Although the pre-trained models have not been fine-tuned on the downstream datasets, the expected risk (test error) on the target task still decreases with the increasing model capacity as shown in Figure 2(a). Formally, when the capacity of pre-trained model $f_{\bm{\hat{\theta}}}$ at step $t+1$ is larger than that at step $t$, we have $\mathcal{R}_{\mathcal{L}}(f_{\bm{\hat{\theta}}(t+1)})<\mathcal{R}_{\mathcal{L}}(f_{\bm{\hat{\theta}}(t)})$. Therefore, the sequence of pre-trained models with different capacities is viable and can be constructed efficiently without fine-tuning. Specifically, we directly produce the classifiers with different capacities by masking the initial pre-training weights with ratios ranging from $\{2\%,4\%,...,98\%,100\%\}$. Note that parameters with small L1 norm are masked first (Han et al., 2015). We provide a concrete formulation of the pre-training weights masking in Appendix B. To calculate the loss integral, i.e., Learning Complexity, the average category feature acts as the class prototype, and we further adopt the Monte Carlo method to approximate the loss integral (Caflisch, 1998) for efficiency, instead of enumerating all models with different capacities:
@@
To empirically verify the asymptotic convergence and decreasing generalization error, we visualize the test loss trend in Figure 2(b). Obviously, $f_{\bm{\theta}(0)}$ gradually converges to $f_{\bm{{\hat{\theta}}}}$ with better generalization performance. Following the Definition 1, the optimization trajectory of SGD during the fine-tuning is a viable learning path, and we calculate the integral of classification loss over the fine-tuning process as the optimization-based learning complexity.
@@ 3.2 FlexRand: Flexible Under-sampling with Randomness
From our previous analysis, we show that DLC can discriminate the hard samples from the easy. To identify informative samples with DLC, we further design a flexible under-sampling strategy with randomness, named FlexRand, replacing the common top-K strategy. For the multinomial distribution of under-sampling, FlexRand aims to flexibly adjust sampling preference at different data regimes. On the other hand, the subset from FlexRand should not be significantly shifted away from the original data distribution. To achieve these objectives, we randomly select the same number of samples from the easy and hard intervals respectively, whose lengths are flexibly tuned with a splitting hyperparameter $\gamma$ to accommodate different data regimes. The sampling probability for $(\bm{x},y)$ can be formulated as follows:
@@
In a data-poor regime, easy samples are more informative, and the opposite holds when more samples can be preserved for fine-tuning (Sorscher et al., 2022). However, the top-K (Easy or Hard) strategy, denoted as T-E or T-H, fixes sampling preference across different data regimes and shows inferior downstream performance compared to the random as shown in Figure 4(b) & 4(d). On the contrary, FlexRand can flexibly adjust sampling preference with a splitting hyperparameter $\gamma$ to accommodate different data regimes.
@@
Figure 4: Harmful distribution shift from the top-K under-sampling strategies. We conduct downstream dataset pruning with four under-sampling strategies according to the optimization-based learning complexity and DLC. (a): Distribution shift comparison of four subsets for optimization-based learning complexity. (b): Downstream performance comparison of four subsets for optimization-based learning complexity. (c): Distribution shift comparison of four subsets for DLC. (d): Downstream performance comparison of four subsets for DLC.
@@ 3.3 Extension to Instruction Dataset Pruning
Recently, large language models (LLMs) (Touvron et al., 2023a; b; Jiang et al., 2024) driven by the PT-FT paradigm have shown incredible capabilities across a wide range of language tasks. Despite the availability of pre-trained weights, the ever-increasing number of parameters still incur heavy computation overhead for the fine-tuning, which involves instruction-tuning (Ouyang et al., 2022) and reinforcement learning with human feedback (RLHF) (Knox & Stone, 2011) to align the pre-trained LLMs with human preferences. On the other hand, recent works (Zhou et al., 2023a; Cao et al., 2023) indicate that almost all knowledge in LLMs is learned during pre-training, and only limited instruction tuning data is necessary to teach models to produce high-quality output. Therefore, we further extend the learning complexity to prune instruction datasets for efficient large language models fine-tuning (Dubey et al., 2024; Team et al., 2024). Specifically, we construct the learning path by masking different numbers of weights in the pre-trained large language models (LLMs), which are sorted by L1 norm in ascending order. Due to the prediction difference in image classification and text generation, we replace the original loss function $\mathcal{L}(f(\bm{x};\bm{\theta},y)$ with:
@@ 4 Experiments
In this section, we present the diverse downstream image and instruction datasets pruning benchmarks and empirically validate the effectiveness and efficiency of the proposed method. Moreover, we perform ablation studies to understand better how different components and hyperparameters influence the performance. The code is available in the supplementary material.
Table 1: Average classification accuracy over 5 diverse downstream datasets and pruning ratios with varying model architectures. Bold numbers are the optimal results, and underline numbers are suboptimal results. Detailed results can be found in Appendix C.2.2.
@@
Figure 5: Trends of average accuracy gap over the random method at different pruning ratios. Detailed results can be found in Appendix C.2.2.
@@ 4.1.1 Downstream Image Dataset Pruning Benchmark
For a comprehensive evaluation, we adopt pre-training encoders with different depths, architectures, and pre-training paradigms. In detail, the pre-trained encoders consist of ResNet-18, ResNet-50 (He et al., 2016), ViT-Small, and ViT-Base (Dosovitskiy, 2020). Those models are pre-trained on the canonical ImageNet-1K (Deng et al., 2009) dataset via fully and weakly supervised learning (Yalniz et al., 2019) respectively.
@@
Generally, the downstream dataset is domain-specific and different from the pre-training dataset ImageNet-1K. Therefore, we choose diverse downstream datasets from 5 domains (Islam et al., 2021) to construct the large-scale benchmark, including CXRB10222CXRB10 is created by selecting 10 balanced classes from the ChestX-ray14 (Borghesi & Maroldi, 2020)., DeepWeeds (Olsen et al., 2019), DTD (Cimpoi et al., 2014), FGVCAircraft (Maji et al., 2013), and Sketch (Eitz et al., 2012). For hyperparameter tuning, we split 20% as the validation set.
@@ 4.1.2 Downstream Instruction Dataset Pruning Benchmark
To align the based pre-trained models, we adopt 2 different instruction fine-tuning datasets: Alpaca Cleaned (Taori et al., 2023) and Dolly & HH-RLHF [^3] https://huggingface.co/datasets/mosaicml/dolly_hhrlhf. More setup details are presented in Appendix C.1.
@@ 4.2.1 Fine-tuning for Image Classification
Our method achieves superior accuracy with much less time costs.
@@
In addition to the accuracy evaluated on different models and downstream setups, we also record the time cost of each baseline, including hyperparameter tuning associated with the scoring estimation and under-sampling. In Table 1, time results averaged over 5 downstream datasets are presented. Benefiting from the training-free scoring function and efficient hyperparameter tuning, the proposed method significantly reduces the pruning cost by 35$\times$ compared with the one-shot fine-tuning baselines while achieving state-of-the-art downstream performance. Notably the acceleration is more remarkable with ever-increasing model capacities. Therefore, our method gives rise to the efficient fine-tuning of the over-parameterized pre-trained models.
Our method works well at various pruning ratios.
For efficacy verification of our method on diverse downstream setups, we compare the accuracy gap trends over the random strategy for each fine-tuning dataset. As shown in Figure 5, the proposed method shows consistent improvement for each downstream dataset at different data regimes, indicating the efficacy of the flexible under-sampling strategy with randomness. Overall, the advantage is coherent and more evident when fine-tuning with fewer samples. For example, our method outperforms the random baseline by 2.6% when removing 90% samples from the Sketch dataset. Despite the randomness in the custom under-sampling, our method demonstrates stability with smaller performance variance. However, the compared methods show inferior performance with fluctuating trends, which reveals the drawbacks of the fixed sampling principle and the challenge of diverse downstream dataset pruning benchmarks.
@@ 4.2.2 Fine-tuning LLMs for Text Classification
Figure 6: Results for different ablation studies. (a): Ablation on the learning path size. (b): Ablation on the masking strategies. (c): Ablation on the under-sampling strategies. T-H, T-E, and T-F denote the top-K Hard, Easy, and Flexible under-sampling, respectively. (d): Ablation on the splitting $\gamma$. Blue, green, and orange dotted lines correspond to 10%, 20%, and 30% of the data, respectively.
@@ 4.3 Ablation Studies
To better understand why our method can obtain superior efficiency and efficacy on the downstream dataset pruning task, we perform extensive ablation studies on the learning complexity metric, under-sampling strategy, and corresponding hyperparameters. We also present a comparison of FlexRand and Dataset Quantization (Zhou et al., 2023b) in Appendix C.2.4. As for the setups, we fine-tune the fully pre-trained ResNet-18 on 5 downstream datasets with 9 pruning ratios with default parameters.
@@
In Figure 6(a), we investigate the effect of the learning path size on the ranking correlation and downstream accuracy. In particular, we calculate the ranking correlation by the class-averaged Spearman correlation coefficients between the optimization-based learning complexity and masking-based. The results show the ranking correlation between the two scoring functions is significant and saturated at size 5. As a result, the proposed method can estimate the learning complexity efficiently and establish superior classification performance.
@@
To examine the effects of interval splitting, we set the $\gamma$ by traversing {0.05, 0.10, …, 0.95}, and present the results in Figure 6(d). We can find that small splitting values ($\gamma<0.5$) generally deliver better performance in data-poor regimes. Therefore, our method outperforms the random baseline. While $\gamma$ tuned by the linear classifier does not always achieve the optimal, it significantly reduces the cost of hyperparameter tuning and establishes superior performance.